OpenStudy (anonymous):

Simplify: square root of 2 over cube root of 2 (It won't let me write it with the radical symbol) Which one is the answer? 2 to the power of 1 over 6 2 to the power of 1 over 3 2 to the power of 5 over 6 2 to the power of 3 over 2

OpenStudy (anonymous):

Also, how would I graph y = the square root of the quantity x plus 6?.

OpenStudy (anonymous):

@ranga Could you help? I'm not a creeper or anything it's just I've seen you answer other people's questions and the answers were very helpful.

OpenStudy (ranga):

\[\Large \frac{ \sqrt{2} }{ \sqrt[3]{2} } = \frac{ 2^{1/2} }{2^{1/3} } = 2^{1/2-1/3} = 2^{1/6} \]

OpenStudy (anonymous):

Thank you so much! For the second question, where would the line begin? The options are it beginning at: -8 -7 -6 -4

OpenStudy (ranga):

Is the graph for \[y = \sqrt{x+6} \quad or \quad y = \sqrt{x} + 6 \]

OpenStudy (anonymous):

the first one

OpenStudy (ranga):

You cannot take the square root of a negative number. square root of 0 and positive numbers are okay but no negative numbers. y = sqrt(x + 6) x + 6 >= 0 x >= -6 So the graph will satrt at x = -6. That is the lowest value allowed for x.

OpenStudy (anonymous):

Thats what I thought :) Can you help me with one more question?

OpenStudy (ranga):

I can try one more here.

OpenStudy (anonymous):

What is 2 to the power of 3 over 2 equal to? square root of 8 cube root of 8 cube root of 16 square root of 16

OpenStudy (ranga):

\[\Large 2^{3/2} = (2^3)^{1/2} = 8^{1/2} = \sqrt{8}\]

OpenStudy (anonymous):

One second

OpenStudy (anonymous):

Would it look like 2^~/8 ^ that is supposed to symbolize the radical.

OpenStudy (ranga):

The answer I have above is: radical 8 or square root of 8.

OpenStudy (ranga):

Not sure what you are asking.